OpenStudy (anonymous):

How do I find these answers I aways get confused.. How do I know if is is a solution There are infinitely many solutions There is no solution

OpenStudy (anonymous):

an example would be 8a+9b=3 && 8a+b=27 Idk what I would ddo

OpenStudy (anonymous):

Isolate one variable then plug it into other sides and solve. Let me show you how to solve it. Let isolate b in second equation because it's easy. 8a + b = 27 b = 27 - 8a now plug it into other equation. 8a + 9b = 3 8a + 9(27 - 8a) = 3 Now solve for a. 8a + 9(27) - 9(8a) = 3 8a + 243 - 72a = 3 Subtract both sides by 243: 8a - 72a = 3 - 243 8a - 72a = -240 Combine the like terms. (8 - 72)a = -240 -64a = -240 Divide both siddes by -64. a = -240/-64 a = 240/64 a = 15/4 Now plug the value of a into other equation and find b. 8a + b=27 8(15/4) + b=27 30 + b = 27 Subtract both sides by 30: b = 27 - 30 b = -3 So a = 15/4 and b = -3 Hope this helps.

OpenStudy (anonymous):

so what would the ordered pay be???? (15,4)

OpenStudy (anonymous):

This is called checking the consistency of linear equations, you will get this by practice when you solve many equations for now for identification: if two equations are given : a1x+b1y+c=0 & a2x+b2y+c=0 then if \[a1/a2 \neq b1/b2\] then one unique solution, if \[a1/a2=b1/b2=c1/c2 \] then infinite solution and if \[a1/a2=b1/b2 \neq c1/c2\] then it is no solution. I hope this will help....moreover you can refer this: http://ncert.nic.in/ncerts/textbook/textbook.htm?jemh1=3-14 or just go to the site: http://ncert.nic.in/ncerts/textbook/textbook.htm on the upper tab put class X, Subject Mathematics and book title Mathematics.....Open Chapter 3,read through you will be able to grasp.

OpenStudy (anonymous):

@Talyababyyy hope you got it......check on the site you will understand it really well..... http://ncert.nic.in/ncerts/textbook/textbook.htm?jemh1=3-14